# Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect

### v1

@@ Abstract
Conclusion.– Within this simple model, we observed that the Unruh effect can be detected with much lower accelerations using the PBP in the presence of reflecting boundaries. With a single mirror, the effect can be observed with acceleration as low as $\sim 10^{10}m/s^{2}$, with a significant separation between the atom and the boundary. With shorter separations, one can still detect the phase with accelerations of a few orders less than that of a free space. With double-mirrors, the effect is more apparent and situation is much more inspiring to build experimental setup. Here, one needs acceleration to detect the phase can be as low as $\sim 10^{9}m/s^{2}$ depending on the separation between the boundaries and the atom’s position inside them for energy gap $\sim$ GHz. The order of the required acceleration can be $10^{3}$ times less if one can utilize a qubit with the energy gap in MHz order. Also, further reducing the acceleration is possible by exploring other possible parameter ranges with enhanced numerical capabilities.
Our study shows that boundaries can significantly influence the geometric phase with much lower accelerations. In fact the minimum observable phase can be detected with very low acceleration when two mirrors are kept within a finite distance. The required acceleration is very low compared to the earlier proposals and hence can be a very strong candidate for construction of an apparatus to detect Unruh effect. However, looking at the strong effects of the mirrors, we may expect cavity arrangement may show further improvement in the PBP. Moreover here we did not incorporate the effects of edges of the mirrors. Incorporation of all these may be very important in this context. We keep these investigation for future.

### v2

@@ Abstract
Conclusion.– Within this simple model, we observed that the Unruh effect may be detected with much lower accelerations using the PBP in the presence of reflecting boundaries. All the parameters used here are close to experimentally feasible ranges, and shows significant influence of mirrors on low-acceleration PBP. With a single mirror, the effect can be observed with acceleration as low as $\sim 10^{10}m/s^{2}$, with a significant separation between the atom and the boundary. With shorter separations, one can still detect the phase with accelerations of a few orders less than that of a free space. With double-mirrors, the effect is more apparent and situation is much more inspiring to build experimental setup. Here, one needs acceleration to detect the phase can be as low as $\sim 10^{9}m/s^{2}$ depending on the separation between the boundaries and the atom’s position inside them for energy gap $\sim$ GHz. In earlier studies without presence of mirrors, the minimum acceleration value to detect the phase difference was $\sim 10^{17}m/s^{2}$, showing the magnificent influence of the mirrors. The order of the required acceleration can be $10$ times less if one can utilise a qubit with the energy gap in 100 MHz order. Also, further reducing the acceleration is possible by exploring other possible parameter ranges with enhanced numerical capabilities.
Our study shows that boundaries can significantly influence the geometric phase with much lower accelerations. In fact the minimum observable phase can be detected with very low acceleration when two mirrors are kept within a finite distance. The required acceleration is very low compared to the earlier proposals and hence can be a very strong candidate for construction of an apparatus to detect Unruh effect. Consequently we propose a schematic idea of experimental setup to witness such effect. In this connection we should be careful about the choice of substance to provide acceleration through the thermal gradient technique. This is because the interaction between the substance and the atom and its temperature can affect the coherence nature of our detector atom. Providing acceleration by the thermal gradient technique usually done by the substances, like Na-K, Sn–Zn, Sn–Pb, Sn–Cd, etc (see, [39]). However it must be ensured that the interaction between the fundamental particle of the substance and our detector atom must be negligible such that it can escape from the substance. Therefore when the atom leaves it, the acceleration is equal to the desired acceleration ($\sim 10^{9}m/s^{2}$).
Here one should note that the width of this substance is extremely small ($\Delta{x}\sim 10^{-5}m$, as mentioned in the earlier section). Hence, the time spend in the substance by the accelerated atom is extremely small. Also one should note that the substance does not directly influence the atom, it can only influence the atom through its temperature. Since $A$ is calculated in $\lambda^{2}$-order, one has $e^{-4A\tau}\approx 1$. In this case if the state of the atom before entering into the substance is given by |$\ket{\psi(0)}$, then due to the temperature effect, the state will remain in the same form with $\theta\to\theta^{\prime}=\theta+C$, where $C$ is a constant. This can be obtained from the analysis of [26] (particularly see Eq. (10) and (11) of [26]). Thus our analysis will remain valid with $\theta$ is replaced by $\theta^{\prime}$ and then by choosing $\theta^{\prime}=\pi/4$. Similar change will also happen for inertial atom. Therefore one have to prepare the initial states accordingly.
However, looking at the strong effects of the mirrors, we may expect cavity arrangement may show further improvement in the PBP. Moreover here we did not incorporate the effects of edges of the mirrors. Also correction due finite temperature is neglected. The interaction in our study can be considered as between one of the components of the electromagnetic field and dipole. Then for practical purpose a derivative coupling has to be considered. However, as discussed in [23], we expect that the order of acceleration will not change. Incorporation of all these may be very important in this context. We keep these investigation for future.
@@ Acknowledgements.
Appendix
Appendix A Convergence with $\text{max}|n|$
In the manuscript we have always taken the maximum values of $n$ upto some particular value. Here we provide the justification for same. We will see that $\delta\phi_{B}$ will remain constant after the particular values we have chosen. In Fig. 4 we have plotted $\delta\phi_{B}$ with respect to max$|n|$ and we show that $\delta\phi_{B}$ does not change after the particular values. Therefore after these max$|n|$ values the contributions are negligible.
[…]
In Fig. 4(a) we can see that for $a/\omega_{0}c=10^{-5}$ the $\delta\phi_{B}$ becomes constant after max $|n|=0.1\,million$. For $a/\omega_{0}c=10^{-6}$ and $10^{-7}$ the quantity $\delta\phi_{B}$ does not change after max $|n|=0.4\,million$. In Fig. 4(b) the quantity $\delta\phi_{B}$ for $L\omega_{0}/c=500$ and $1000$ remain constant after max $|n|=0.5\,million$. This justifies our particular chosen values in our manuscript.